# Relating Checkpoint Update Probabilities to Momentum Parameters in Single-Loop Variance Reduction Methods

### v1

@@ Abstract
Variance reduction (VR) is a crucial tool for solving finite-sum optimization problems, including the composite general convex setting, which is the focus of this work. On the one hand, denoting the number of component functions as $n$ and the target accuracy as $\epsilon$, some VR methods achieve the near-optimal complexity $\widetilde{\mathcal{O}}\left(n+\sqrt{n}/\sqrt{\epsilon}\right)$, but they all have complicated nested structure and fail to provide convergence guarantee for the iterate sequence itself. On the other hand, single-loop VR methods, being free from the aforementioned disadvantages, have complexity no better than $\mathcal{O}\left(n+n/\sqrt{\epsilon}\right)$ which is the complexity of the deterministic method FISTA, thus leaving a critical gap unaddressed. In this work, we propose the Harmonia technique which relates full-gradient point (i.e., checkpoint) update probabilities to momentum parameters in single-loop VR methods. Based on this technique, we further propose to vary the growth rate of the momentum parameter, creating a novel continuous trade-off between acceleration and variance reduction, controlled by the key parameter $\alpha\in[0,1]$. The proposed techniques lead to following favourable consequences. First, several known complexity of quite different algorithms are re-discovered under the proposed unifying algorithmic framework Katyusha-H. Second, under an extra mild condition, Katyusha-H achieves the near-optimal complexity for $\alpha$ belonging to a certain interval, highlighting the effectiveness of the acceleration-variance reduction trade-off. Moreover, by choosing smaller $\alpha$, Katyusha-H can achieve the same complexity (in the big-$\mathcal{O}$ sense) with less per-iteration cost. Last, without extra conditions, Katyusha-H achieves the complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ with $\alpha=1$ and proper mini-batch sizes. The proposed idea and techniques may be of general interest beyond the considered problem in this work.
@@
finite-sum optimization, single-loop methods, variance reduction, acceleration, Katyusha momentum
@@ 1 Introduction
where $f(x):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)$ with $f_{i}(x)$ being convex and smooth, and $l(x)$ is convex, lower semicontinuous (but possibly non-differentiable) function which admits an efficient proximal operator. Problems of this type are ubiquitous in statistics and machine learning such as regularized empirical risk minimization [1].
When $n$ is large, the computation of $\nabla f(x)$ is expensive since it requires computing $n$ component gradients. Gradient-based methods (e.g., the forward-backward splitting algorithm [2] and its inertial version [3, 4]) require access to the full gradient at every iteration, hence are unpractical. By utilizing the finite-sum structure of (1), stochastic gradient methods use the gradient of one or a few component(s) $\nabla f_{i}$ to construct the stochastic gradient estimator. For example, stochastic gradient descent (SGD) uses $\nabla f_{i}(x)$ (ii is chosen from $n]:=\{1,2,\cdots,n\}$ according to the discrete uniform distribution) as an unbiased estimator of $\nabla f(x)$ in each iteration, hence enjoys a low per-iteration cost. However, due to the non-diminishing variance of its gradient estimator, the convergence speed of SGD is very slow. Even when the objective is smooth and strongly convex, SGD can only achieve sublinear convergence.
Variance reduction (VR) methods [5, 6, 7, 8, 9, 10, 11, 12] utilize history information to construct variance-vanishing stochastic gradient estimators, enjoying both low per-iteration cost and fast convergence speed. In fact, the convergence rate of VR methods improves significantly upon SGD, and matches that of traditional gradient-based methods. For example, VR methods achieve the linear convergence rate in the smooth strongly convex problems.
@@
where $i_{t}$ is chosen from $[n]$ according to the discrete uniform distribution at $t$th iteration. Then the iterate is iterated with the formula $x_{t+1}=x_{t}-\eta\tilde{\nabla}f_{i_{t}}(x_{t})$, where $\eta>0$ is the step size. After $m$ iterations, the checkpoint is updated (e.g., it can be chosen as the average of last $m$ iterates) and the aforementioned procedure is repeated. Hence the vanilla SVRG method, along with many other VR methods [13, 14, 7, 15], have nested algorithmic structure.
@@
VR methods achieve well-known success in solving the composite general convex problem (1) under the incremental first-order oracle (IFO) framework [16], which evaluates the efficiency of an algorithm by the number of evaluating component gradients $\nabla f_{i}(x)$ needed to achieve an $\epsilon$-accurate solution. The vanilla SVRG relies on the strong convexity of $f$ to combine different outer iterations in the proof, hence cannot directly solve general convex problem theoretically. SVRG++ [17] resolves this issue by introducing the technique of doubling the inner loop length (i.e., $m_{s}=2^{s}\cdot m_{0}$ where $m_{s}$ is the inner loop length of the $s$th outer iteration), achieving the IFO complexity $\mathcal{O}(n\log(1/\epsilon)+1/\epsilon)$.
@@
Katyushans [8] uses the Katyusha momentum [8] to mitigate possible negative effects of Nesterov acceleration in stochastic setting. It achieves the complexity $\mathcal{O}(n+n/\sqrt{\epsilon})$, and the Katyusha momentum turned out to be instrumental for the acceleration of VR methods in the composite general convex problem. MiGnsc[18] simplifies the iteration structure of Katyushans and achieves essentially the same complexity $\mathcal{O}(n+n/\sqrt{\epsilon})$. Using the Katyusha momentum in inner iterations, ASVRG [19] achieves the complexity $\mathcal{O}(n\log(1/\epsilon))+\sqrt{n}/\sqrt{\epsilon})$, which nearly matches the theoretical lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ [20]. By utilization of the Katyusha momentum, the acceleration techniques in [21, 22], and a multi-stage inner loop length selection scheme, Varag [23] also achieves the complexity $\mathcal{O}(n\log(n)+\sqrt{n}/\sqrt{\epsilon})$. VRADA [24] incorporates the Katyusha momentum with a novel initialization technique, achieving the complexity $\mathcal{O}(n\log\log(n)+\sqrt{n}/\sqrt{\epsilon})$. Based on these methods, some researchers employ the AdaGrad idea [25] to compute step sizes [26], maintaining similar complexity under additional assumptions such as the compactness of the underlying set which the optimization variable belongs to.
@@
As introduced above, some VR methods can achieve the near-optimal complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$, where the $\widetilde{\mathcal{O}}$ notation hides logarithmic factors. However, all such methods have the nested algorithmic structure, which brings two major drawbacks. First, this complicated structure renders the implementation of the algorithm more troublesome. Second, these methods only has convergence guarantee for the checkpoint sequence. The checkpoint is usually not the last iterate of the last $m$ inner iterations [^1] Theoretically, choosing the last iterate of last $m$ inner iterations to be the new checkpoint deteriorates the complexity. Hence existing methods usually choose the checkpoint with certain (weighted) average, although this is empirically undesirable. and a new checkpoint can be obtained only after an entire outer iteration. For similar reasons, indirect methods [27], which adds strongly convex regularization terms to (1) with decreasing regularization parameters and uses VR methods to solve resultant strongly convex problems, are also undesirable.
@@
In contrast, single-loop VR methods are easy to implement and typically provide convergence guarantee for the iterate sequence itself. However, the literature on single-loop VR methods for the composite general convex problem (1) is relatively limited. The complexity of single-loop VR methods are no better than $\mathcal{O}(n+n/\sqrt{\epsilon})$, which is exactly the complexity of FISTA [4]. Note that FISTA is an extension of Nesterov’s accelerated gradient method [28] for the composite optimization problem, and it requires full gradient computation in every iteration, hence is a gradient-based method.
@@
In this line of works, L-SVRG [9] and L-Katyusha [9] are first single-loop VR methods. In L-SVRG, the checkpoint, denoted as $w_{t}$ (where $t$ is the iteration number), is updated in a randomized way. Concretely, at the end of the $t$th iteration, $w_{t+1}$ is replaced by the new iterate [^2] Once the replacement occurs, a full gradient is computed at $w_{t+1}$. with probability $p$ or remains $w_{t}$ otherwise, where $p\in[0,1]$ is a constant which takes a small value, e.g., $1/n$. However, L-SVRG and L-Katyusha require the smooth part $f$ to be strongly convex and the non-smooth part $l\equiv 0$, hence cannot solve the composite general convex problem (1). L2S [29] is a single-loop VR method that uses the SARAH gradient estimator [7] and a constant checkpoint update probability, having the complexity $\mathcal{O}(n+\sqrt{n}/\epsilon)$. However, it requires that $l\equiv 0$ in (1) and the convergence guarantee is only for the expected squared gradient norm, not the expected functional gap $\mathbb{E}[F(\cdot)-\inf F]$ which is usually adopted.
@@
Without using the Katyusha momentum, Driggs et al. [30] proposed a universal single-loop acceleration framework for several frequently used variance reduced gradient estimators based on the Nesterov acceleration scheme. Under the composite general convex setting, their method has the convergence rate $\mathcal{O}(c_{n}/T^{2})$, where $T$ is the number of total iterations and $c_{n}$ is an $n$-dependent constant. Curiously, using mini-batch cannot improve the complexity of this scheme, regardless of which estimator is employed. For example, when the SVRG estimator is used, as the mini-batch size $b$ grows, $c_{n}$ is decreased but the per-iteration cost is increased. As a result, the gain in the iteration complexity and the loss in the per-iteration cost cancels perfectly, leading to the complexity $\mathcal{O}(n+n/\sqrt{\epsilon})$ for any $b\in[n]$. The checkpoint update probability in [30] is also a constant $p=b/n$.
@@
Note that setting $p$ to a constant is essential for the proof of [30], since Driggs et al. estimate the summation of the bias term and mean squared error of the gradient estimator respectively along the entire optimization process. The constant probability setting renders certain coefficient (in the proof) essentially the truncation of a geometric series hence upper-bounded and independent of the total iteration number $T$. If a non-constant probability setting is to be directly applied, a positive lower bound of the probabilities is likely to be required to prevent aforementioned coefficients from exploding. But in this way, the complexity arguably relates to the aforementioned lower bound in essential, hence cannot be improved in terms of the $n$-dependence.
@@
After the completion of this manuscript, we came to realize the article [31] in which the SIFAR algorithm is proposed [^3] During the literature research, we thoroughly searched for all related articles of [24], which proposes the VRADA method enjoying the currently best complexity among nested VR methods, and [30], which is the only known single-loop VR method with momentum technique solving (1) to the best of our knowledge. The reason that we did not realize [31] is partly due to that [31] accidentally missed these articles.. The SIFAR algorithm is a single-loop VR method with acceleration technique and the checkpoint update probability $p_{t}$ varies between different stages. [31] denotes the iteration in which the checkpoint is updated for the first time as $t_{1}$, which follows the geometric distribution. Then there are three stages of SIFAR: $p_{t}\equiv 1/(n+1)$ for the first stage $0\leq t\leq t_{1}$, $p_{t}=4/(t-t_{1}+3\sqrt{n})$ for the second stage $t_{1}<t\leq t_{1}+n+3-3\sqrt{n}$, and $p_{t}\equiv 4/(n+3)$ for the third stage $t>t_{1}+n+3-3\sqrt{n}$.
The complexity $\mathcal{O}(n\min\{1+\log(1/(\epsilon\sqrt{n})),\log\sqrt{n}\}+\sqrt{(nL)/\epsilon})$ is claimed in [31]. However, we believe that the proofs in [31] are incorrect after careful consideration. First, Theorem $1$ of [31] is incorrect. In Lemma $3$ of [31], an upper bound is given for the quantity $\mathbb{E}\left\{[\eta_{t-1}/(p_{t-1}\theta_{t-1}^{2})][f(w_{t})-f(x^{*})]\right\}$ for any $t>t_{1}$. Since $t_{1}$ is a random variable (r.v.) that may take any positive integer value in the probability space, $t$ should also be a r.v. to satisfy the inequality $t>t_{1}$, hence the coefficient $\eta_{t-1}/(p_{t-1}\theta_{t-1}^{2})$ is a r.v. since it depends on $t$ and $t_{1}$. In [31], this coefficient is directly taken out of the expectation $\mathbb{E}$ and both sides of the inequality are multiplied by $(p_{t-1}\theta_{t-1}^{2})/\eta_{t-1}$ to obtain an upper bound on the concerned quantity $\mathbb{E}(f(w_{t})-f(x^{*}))$. However, this reasoning is illegal and the upper bounding of $\mathbb{E}(f(w_{t})-f(x^{*}))$ is elusive. Hence, Theorem $1$ of [31] is incorrect and the difficulty there is critical.
Consider two correlated r.v.’s $X_{1}$ and $X_{2}$, the estimating of terms of the form $\mathbb{E}(X_{1}X_{2})$ [^4] Sometimes such terms take the form of the expectation of a sum of products of correlated r.v.’s, i.e., $\mathbb{E}\sum_{t=0}^{T}X_{t}^{(1)}X_{t}^{(2)}$, where $X_{t}^{(1)}$ and $X_{t}^{(2)}$ are correlated r.v.’s for $0\leq t\leq T$. But we take the simple $T=0$ case for the ease of discussion. is of fundamental importance in the analysis of randomized methods. In some cases, we know how to estimate $\mathbb{E}(X_{2})$ or $\mathbb{E}(X_{2})$ is the concerned quantity, and we want to apply the known inequality or obtain $\mathbb{E}(X_{2})$ alone. For example, such terms arise when the step size is computed via the current and/or some previous (incremental) gradients, e.g., the step size is determined by the Barzilai-Borwein-type formulae [32] or the AdaGrad-type formulae [25]. Then whether one can successfully estimate such terms determines the success or failure of the convergence analysis [33]. And when estimating them are feasible, the way of doing the estimates critically affects the ultimate theoretical complexity [34]. Hence, the problem that Theorem $1$ of [31] faced is beyond a trivial mis-reasoning.
Second, Corollary $1$ of [31] is incorrect. Due to the fact that $t_{1}$ is a r.v., the iterations needed by the SIFAR algorithm to achieve an $\epsilon$-accurate solution is also a r.v. To the best of our knowledge, this is in contrast to other existing single-loop VR methods, where the required iterations is deterministic given the target accuracy $\epsilon$. Hence the analysis of the complexity should be conducted differently. For a single-loop VR method with probabilistic update of the checkpoint, denote the computational cost of the $t$th iteration as $C_{t}$ (note that $C_{t}$ is a r.v.), and denote the iterations needed to reach an $\epsilon$-accurate solution as $T(\epsilon)$. Then the total cost in expectation is $\mathbb{E}\sum_{t=0}^{T(\epsilon)}C_{t}$. When $T(\epsilon)$ is deterministic, we have $\mathbb{E}\sum_{t=0}^{T(\epsilon)}C_{t}=\sum_{t=0}^{T(\epsilon)}\mathbb{E}(C_{t})$. However, when $T(\epsilon)$ is a r.v., certain propositions, e.g., the Wald’s theorem or its generalized versions [35, 36], should be applied in such case (if they are applicable at all under the design and characteristics of the SIFAR algorithm).
@@
In all, despite the appealing novelty and innovation of [31], we believe that the convergence analysis and complexity analysis concerning the general convex setting are incorrect, hence the corresponding complexity is invalid.
@@
Inspired by the aforementioned theoretical and practical merit of the single-loop methods, we take this line of research further in this work. We propose the Harmonia technique which relates checkpoint update probability $p_{t}$ to momentum parameter $\alpha_{t}$ in single-loop VR methods to construct provably better algorithms. Our main contributions are summarized as follows:
@@
  • Along with the existing Nesterov acceleration framework and the Katyusha momentum, we propose the Harmonia technique to accelerate single-loop VR methods in the composite general convex problem. With mini-batch size properly set, the proposed Katyusha-H method achieves the complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$, which significantly improves upon the previous best $\mathcal{O}(n+n/\sqrt{\epsilon})$ of single-loop VR methods and matches the lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ up to logarithmic factors.
  • Based on the Harmonia technique, we further propose to vary the growth rate of the momentum parameter, creating a novel continuous trade-off between acceleration and variance reduction, controlled by the key parameter $\alpha\in[0,1]$. Under an extra mild condition, Katyusha-H achieves $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ for $\alpha$ belonging to an explicitly given interval. In many cases, $\alpha=1$ which corresponds to the largest allowable acceleration strength does not belong to the aforementioned interval, hence showing the effectiveness of the proposed acceleration-variance reduction trade-off. Moreover, by choosing smaller $\alpha$, Katyusha-H maintains the near-optimal complexity while having less per-iteration cost.
  • Several known complexity of existing algorithms that use different optimization techniques and exhibit different structure are re-discovered under the unifying algorithmic framework Katyusha-H.
  • Based on the conducted theoretical analysis, we provide intuitive interpretations for the proposed $p_{t}-\alpha_{t}$ correspondence and the key parameter $\alpha$, which may be of general interest.
@@
The rest of this article is organized as follows: Section 2 introduces notations, definitions and assumptions. Section 3 presents the proposed Katyusha-H method and related preliminary explanations. Section 4 presents our theoretical findings and detailed discussions and interpretations concerning the theoretical results, where most of the proofs are deferred to the Appendix. Section 5 summarises the proposed idea and technique from a high-level perspective and relist the contributions briefly.

### v2

@@ Abstract
We propose a single-loop variance-reduced acceleration framework, which relates checkpoint update probabilities to momentum parameters, for solving the composite general convex problem where the smooth part has the finite-sum structure. Under the proposed framework, the growth rate of the momentum parameter is further altered, creating a novel continuous trade-off between acceleration and variance reduction, controlled by the key parameter $\alpha\in[0,1]$. A series of novel complexity is obtained, and some complexity of distinct known methods are rediscovered under the unified framework. When the mini-batch size is restricted due to the massive scale of the problem or the computational resource shortage, near-optimal complexity can still be achieved by choosing suitable $\alpha$ for any prefixed target accuracy. Analysis shows that although the considered gradient oracle is exact, acceleration comes with implicit price of heavier variance reduction, hence the obtained optimal $\alpha$ not necessarily corresponds to the largest allowable acceleration strength. Without prefixing the target accuracy, the proposed method achieves the near-optimal complexity $\tilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ to obtain an $\epsilon$-accurate solution under standard assumptions ($n$ is the number of components of the finite-sum), significantly improves upon previous best complexity $\mathcal{O}(n+n/\sqrt{\epsilon})$ of single-loop variance reduction methods, which does not exceed the complexity of the deterministic method FISTA. Numerical experiments demonstrate the efficiency of the proposed method and validate other theoretical findings.
@@
finite-sum optimization, single-loop methods, variance reduction, acceleration
pacs:
Mathematics Subject Classification]90C06, 90C15, 90C25, 68Q25
@@ 1 Introduction
where $f(x):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)$ with $f_{i}(x)$ being convex and smooth, and $l(x)$ is convex, lower semicontinuous (but possibly non-differentiable) function which admits an efficient proximal operator. Problems of this type are ubiquitous in statistics and machine learning such as regularized empirical risk minimization [bottou2018optimization].
When $n$ is large, the computation of $\nabla f(x)$ is expensive since it requires computing $n$ component gradients. Gradient-based methods (e.g., the forward-backward splitting algorithm [lions1979splitting] and its inertial version [nesterov2013gradient, beck2009fast]) require access to the full gradient at every iteration, hence are unpractical. By utilizing the finite-sum structure of (1), stochastic gradient methods use the gradient of one or a few component(s) $\nabla f_{i}$ to construct the stochastic gradient estimator. For example, stochastic gradient descent (SGD) uses $\nabla f_{i}(x)$ (ii is chosen from $n]:=\{1,2,\cdots,n\}$ according to the discrete uniform distribution) as an unbiased estimator of $\nabla f(x)$ in each iteration, hence enjoys a low per-iteration cost. However, due to the non-diminishing variance of its gradient estimator, the convergence speed of SGD is very slow. Even when the objective is smooth and strongly convex, SGD can only achieve sublinear convergence.
Variance reduction (VR) methods [johnson2013accelerating, xiao2014proximal, nguyen2017sarah, allen2018katyusha, kovalev2020don, tan2016barzilai, liu2025improvement, gower2020variance] utilize history information to construct variance-vanishing stochastic gradient estimators, enjoying both low per-iteration cost and fast convergence speed. In fact, the convergence rate of VR methods improves significantly upon SGD, and matches that of traditional gradient-based methods. For example, in contrast to the slow convergence of SGD, VR methods achieve the linear convergence rate in the smooth strongly convex problems.
@@
where $i_{t}$ is chosen from $[n]$ according to the discrete uniform distribution at $t$th iteration. Then the iterate is iterated with the formula $x_{t+1}=x_{t}-\eta\tilde{\nabla}f_{i_{t}}(x_{t})$, where $\eta>0$ is the step size. After $m$ iterations, the checkpoint is updated (e.g., it can be chosen as the average of last $m$ iterates) and the aforementioned procedure is repeated. Hence the vanilla SVRG method, along with many other VR methods [konevcny2017semi, lei2020adaptivity, nguyen2017sarah, horvath2022adaptivity], have nested algorithmic structure.
@@
VR methods achieve well-known success in solving the composite general convex problem (1) under the incremental first-order oracle (IFO) framework [agarwal2015lower], which evaluates the efficiency of an algorithm by the number of component gradient $\nabla f_{i}(x)$ calculations needed to achieve an $\epsilon$-accurate solution. The vanilla SVRG relies on the strong convexity of $f$ to combine different outer iterations in the proof, hence cannot directly solve general convex problem theoretically. SVRG++ [allen2016improved] resolves this issue by introducing the technique of doubling the inner loop length (i.e., $m_{s}=2^{s}\cdot m_{0}$ where $m_{s}$ is the inner loop length of the $s$th outer iteration), achieving the IFO complexity $\mathcal{O}(n\log(1/\epsilon)+1/\epsilon)$.
@@
Katyushans [allen2018katyusha] uses the Katyusha momentum [allen2018katyusha] to mitigate possible negative effects of Nesterov acceleration in stochastic setting. It achieves the complexity $\mathcal{O}(n+n/\sqrt{\epsilon})$, and the Katyusha momentum turned out to be instrumental for the acceleration of VR methods in the composite general convex problem. MiGnsc[zhou2018simple] simplifies the iteration structure of Katyushans and achieves essentially the same complexity $\mathcal{O}(n+n/\sqrt{\epsilon})$. Using the Katyusha momentum in inner iterations, ASVRG [shang2018asvrg] achieves the complexity $\mathcal{O}(n\log(1/\epsilon))+\sqrt{n}/\sqrt{\epsilon})$, which nearly matches the theoretical lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ [woodworth2016tight]. By utilization of the Katyusha momentum, the acceleration techniques in [ghadimi2012optimal, ghadimi2013optimal], and a multi-stage inner loop length selection scheme, Varag [lan2019unified] also achieves the complexity $\mathcal{O}(n\log(n)+\sqrt{n}/\sqrt{\epsilon})$. VRADA [song2020variance] incorporates the Katyusha momentum with a novel initialization technique, achieving the complexity $\mathcal{O}(n\log\log(n)+\sqrt{n}/\sqrt{\epsilon})$. Based on these methods, some researchers employ the AdaGrad idea [duchi2011adaptive] to compute step sizes [dubois2022svrg], maintaining similar complexity under additional assumptions such as the compactness of the underlying set which the optimization variable belongs to.
@@
As introduced above, some VR methods can achieve the near-optimal complexity $\widetilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$, where the $\widetilde{\mathcal{O}}$ notation hides logarithmic factors. However, all such methods have the nested algorithmic structure, which brings two major drawbacks. First, this complicated structure renders the implementation of the algorithm more troublesome. Second, these methods only have convergence guarantee for the checkpoint sequence. The checkpoint is usually not the last iterate of the last $m$ inner iterations [^1] Theoretically, choosing the last iterate of last $m$ inner iterations to be the new checkpoint deteriorates the complexity. Hence existing methods usually choose the checkpoint with certain (weighted) average, although this is empirically undesirable. and a new checkpoint can be obtained only after an entire outer iteration. Since the inner loop length $m$ can be very large, e.g., $m=\mathcal{O}(n)$, the limited convergence guarantee can be troublesome in large-scale problems.
@@
In contrast, single-loop VR methods are easy to implement and typically provide convergence guarantee for the iterate sequence itself. However, the literature on single-loop VR methods for the composite general convex problem (1) is relatively limited. The complexity of single-loop VR methods are no better than $\mathcal{O}(n+n/\sqrt{\epsilon})$, which is exactly the complexity of FISTA [beck2009fast]. Note that FISTA is an extension of Nesterov’s accelerated gradient method [nesterov1983method] for the composite optimization problem, and it requires full gradient computation in every iteration, hence is a gradient-based method.
@@
In this line of works, L-SVRG [kovalev2020don] and L-Katyusha [kovalev2020don] are first single-loop VR methods. In L-SVRG, the checkpoint, denoted as $w_{t}$ (where $t$ is the iteration number), is updated in a randomized way. Concretely, at the end of the $t$th iteration, $w_{t+1}$ is replaced by the new iterate [^2] Once the replacement occurs, a full gradient is computed at $w_{t+1}$. with probability $p$ or remains $w_{t}$ otherwise, where $p\in[0,1]$ is a constant which takes a small value, e.g., $1/n$. However, L-SVRG and L-Katyusha require the smooth part $f$ to be strongly convex and the non-smooth part $l\equiv 0$, hence cannot solve the composite general convex problem (1). L2S [li2020convergence] is a single-loop VR method that uses the SARAH gradient estimator [nguyen2017sarah] and a constant checkpoint update probability, having the complexity $\mathcal{O}(n+\sqrt{n}/\epsilon)$. However, it requires that $l\equiv 0$ in (1) and the convergence guarantee is only for the expected squared gradient norm, not the expected functional gap $\mathbb{E}[F(\cdot)-\inf F]$ which is usually adopted.
@@
Without using the Katyusha momentum, Driggs et al. [driggs2022accelerating] proposed a universal single-loop acceleration framework for several frequently used variance reduced gradient estimators based on the Nesterov acceleration scheme. Under the composite general convex setting, their method has the convergence rate $\mathcal{O}(c_{n}/T^{2})$, where $T$ is the number of total iterations and $c_{n}$ is an $n$-dependent constant. Curiously, using mini-batch cannot improve the complexity of this scheme, regardless of which estimator is employed. For example, when the SVRG estimator is used, as the mini-batch size $b$ grows, $c_{n}$ is decreased but the per-iteration cost is increased. As a result, the gain in the iteration complexity and the loss in the per-iteration cost cancels perfectly, leading to the complexity $\mathcal{O}(n+n/\sqrt{\epsilon})$ for any $b\in[n]$. The checkpoint update probability in [driggs2022accelerating] is also a constant $p=b/n$.
@@
Note that setting $p$ to a constant is essential for the proof of [driggs2022accelerating], since Driggs et al. estimate the summation of the bias term and mean squared error of the gradient estimator respectively along the entire optimization process. The constant probability setting renders certain coefficient (in the proof) essentially the truncation of a geometric series hence upper-bounded and independent of the total iteration number $T$. If a non-constant probability setting is to be directly applied, a positive lower bound of the probabilities is likely to be required to prevent aforementioned coefficients from exploding. But in this way, the complexity arguably relies on the aforementioned lower bound in essential, hence cannot be improved in terms of the $n$-dependence.
@@
After the completion of this manuscript (except for the experimental parts), we came to realize the article [li2025sifar] in which the SIFAR algorithm is proposed. The SIFAR algorithm is a single-loop VR method with acceleration technique and the checkpoint update probability $p_{t}$ varies between different stages. [li2025sifar] denotes the iteration in which the checkpoint is updated for the first time as $t_{1}$, which follows the geometric distribution. There are three stages of SIFAR: $p_{t}\equiv 1/(n+1)$ for the first stage $0\leq t\leq t_{1}$, $p_{t}=4/(t-t_{1}+3\sqrt{n})$ for the second stage $t_{1}<t\leq t_{1}+n+3-3\sqrt{n}$, and $p_{t}\equiv 4/(n+3)$ for the third stage $t>t_{1}+n+3-3\sqrt{n}$.
@@
The complexity $\mathcal{O}(n\min\{1+\log(1/(\epsilon\sqrt{n})),\log\sqrt{n}\}+\sqrt{n/\epsilon})$ is claimed in [li2025sifar]. However, despite the appealing novelty and innovation of [li2025sifar], after careful consideration, we believe that the convergence analysis and complexity analysis concerning the general convex setting are incorrect (detailed discussions are deferred to Appendix E), and the corresponding convergence property and complexity are invalid.
@@
Inspired by the aforementioned theoretical and practical merit of the single-loop methods, we take this line of research further in this work. Our main contributions are summarized as follows:
  • We propose a single-loop variance-reduced acceleration framework, where the checkpoint update probabilities are related to the momentum parameters. By further alternating the growth rate of the momentum parameters, we discover a novel continuous trade-off between acceleration and variance reduction, controlled by the key parameter $\alpha\in[0,1]$. A series of novel complexity are obtained, and some known complexity of methods using different optimization techniques are rediscovered under the proposed unified framework.
  • When the mini-batch size is restricted, the proposed method still achieves near-optimal complexity with proper $\alpha$ choice for any prefixed target accuracy. Analysis reveals that even in the noise-absent case, acceleration comes with implicit price of heavier variance reduction, hence the proposed acceleration-variance reduction trade-off should be considered and can be beneficial in certain scenarios.
@@
  • Setting $\alpha=1$ and the mini-batch size $b=\mathcal{O}(\sqrt{n})$, we prove that the proposed method achieves the complexity $\tilde{\mathcal{O}}(n+\sqrt{n}/\sqrt{\epsilon})$ under standard assumptions. The obtained complexity significantly improves upon previous best complexity $\mathcal{O}(n+n/\sqrt{\epsilon})$ of single-loop VR methods, and matches the lower bound $\Omega(n+\sqrt{n}/\sqrt{\epsilon})$ up to a logarithmic factor.
  • Numerical experiments demonstrate the efficiency of the proposed method and validate other theoretical results.
@@
The rest of this article is organized as follows: Section 2 introduces notations, definitions and assumptions. Section 3 presents the proposed framework and some preliminary explanations. Section 4 presents our theoretical findings and detailed discussions and interpretations concerning the theoretical results, and most of the proofs are deferred to the Appendix. Section 5 gives the results of numerical experiments. Finally, Section 6 summarises the proposed idea and technique from a high-level perspective and relist the contributions briefly.